# Learning with Incomplete Context: Linear Contextual Bandits with Pretrained Imputation

### v2

@@ 6 NUMERICAL EXPERIMENTS
In this section, we validate our theory and algorithm with simulations on synthetic data and the real Taobao Ad Display/Click dataset (Alibaba,, 2018).
@@ 6.1 Synthetic Experiments
PULSE-UCB consists of two phases. In pretraining, a context transition model is learned from $N=1000$ historical time series of length $T_{0}=100$ to predict the latent context $W_{t}$. In the online evaluation, the agent runs for $T=1000$ steps. We compare against two benchmarks: (i) OFUL, a naive agent that ignores $W_{t}$ and uses only $\bm{S}_{t}$; (ii) OFUL-Full, an idealized agent with access to the full context $\bm{Y}_{t}=(\bm{S}_{t},W_{t})$. The cumulative regret, averaged over 30 independent trials, is shown in Figure 1. As expected, OFUL-Full achieves the lowest regret since it observes the full context, while OFUL performs worst by ignoring the missing component. In both the linear and nonlinear settings for the missing context, PULSE-UCB performs nearly as well as OFUL-Full, demonstrating the clear benefit of leveraging a predictive model for the unobserved context.
@@ 6.2 Real-World Experiments
To evaluate our method in a practical setting, we use the public Taobao Ad Display/Click dataset (Alibaba,, 2018), which contains 186,730 advertisement display/click records from Taobao.com. Each record includes 83 features describing user and ad attributes such as gender, age, consumption grade, brand, and category. We embed the features into a 32-dimensional space and partition them into 16 observed features ($\bm{S}_{t}$) and 16 unobserved features (${\bm{W}}_{t}$). All algorithms are evaluated on $80\%$ of the data. For PULSE-UCB, we additionally use the remaining $20\%$ for pretraining the context transition model. The action corresponds to selecting an ad (adgroup ID), and the reward is the binary click feedback (1 if clicked, 0 otherwise). Further preprocessing details are deferred to the Appendix.
We compare PULSE-UCB with three baselines: OFUL, which ignores the missing context; OFUL-Full, which has access to the full context; and CLBBF (Kim et al.,, 2023), designed for bandits with stochastically missing features. We compare these algorithms over $T\approx 1.5\times 10^{5}$ steps, with $K=20$ arms per step, averaging results over 5 runs. Figure 2 shows that PULSE-UCB greatly outperforms OFUL, highlighting the benefit of context reconstruction, and also surpasses CLBBF, whose mechanism struggles under structural missingness. Notably, PULSE-UCB achieves performance nearly indistinguishable from the ideal OFUL-Full, indicating that the pretraining step not only imputes the missing context but also produces a feature representation well suited for linear bandit learning.
@@ Appendix A Additional literature review
AI-assisted decision-making. Recently, there has been growing interest in applying AI, including foundation models, to enhance decision-making. For example, Tianhui Cai et al., (2024); Zhang et al., (2025) design Thompson sampling algorithms for online bandits that treat uncertainty as missing future outcomes, imputing them with pretrained generative models to optimize policies. Chen et al., (2021); Janner et al., (2021) recast offline reinforcement learning as sequence modeling over trajectories to improve decisions, while Lin et al., (2023); Lee et al., (2023) study in-context reinforcement learning, showing that supervised pretraining on past trajectories enables models to approximate algorithms such as LinUCB and Thompson sampling with regret guarantees. Applications include LLM-assisted adaptive experimentation for content delivery (Ye et al.,, 2025) and human–AI collaboration in linear bandits with resource constraints for healthcare (Cao et al.,, 2024). Our work focuses on the missing-context issue, specifically in online contextual bandits, and provides near-optimal regret guarantees that guide the use of pretrained models for imputation in this setting.
@@
Imputation in statistics and ML. Imputation has long been a central strategy across statistics and machine learning for handling missing information. A classical example is the EM algorithm (Dempster et al.,, 1977), which provides a likelihood-based framework for parameter estimation with incomplete data and remains highly influential in this area. In causal inference, imputation is widely used for estimating potential outcomes under counterfactual interventions (Little and Rubin,, 2019). From a statistical learning perspective, recent work has incorporated modern machine learning models to deal with missing responses: Xia and Wainwright, (2024) propose surrogate training that leverages helper covariates to impute pseudo-responses for unlabeled data, yielding prediction improvements with excess risk guarantees, while Angelopoulos et al., (2023) demonstrate that combining a small labeled set with imputed outcomes enables valid confidence intervals and hypothesis tests. Our work contributes to this line of research by extending imputation-based methods to sequential decision-making with partially observed contexts, and quantify the impact of imputation quality on online learning performance.
Appendix B A data-driven approach for choosing $D_{t}$
@@
    $\big|\mu_{p}(\bm{s})-\mu_{\widehat{p}}(\bm{s})\big|\;\leq\;\underbrace{\big|\mu_{p}(\bm{s})-\mu_{\widehat{p}_{0}}(\bm{s})\big|}_{\text{controlled by the UCB}}\;+\;\underbrace{\big|\mu_{\widehat{p}_{0}}(\bm{s})-\mu_{\widehat{p}}(\bm{s})\big|}_{\text{fully data-computable}}.$ (B.1)
@@
$\mathcal{E}_{t}(p,\widehat{p};\bm{s})\;\leq\;\sup_{\bm{s}\in\mathcal{G}_{t}}r_{1-\alpha}(\bm{s})\;+\;\big|\mu_{\widehat{p}_{0}}(\bm{s})-\mu_{\widehat{p}}(\bm{s})\big|,$ (B.2)
@@
    UCB machinery for the reference fit $\widehat{p}_{0}$. Two widely used choices are: (i) local-polynomial regression with robust bias correction (RBC), whose studentized process admits valid simultaneous bands and is robust to MSE-optimal bandwidth choice; the quantiles are obtained via Gaussian/multiplier bootstrap of the sup-statistic; (ii) penalized splines with simultaneous bands via volume-of-tube or bootstrap calibrations. [^2] See, e.g., Calonico et al., (2018, 2022) for RBC-based bands and Krivobokova et al., (2010) for spline bands; multiplier bootstrap for suprema is treated in Chernozhukov et al., (2014).
@@
When the observed context $\bm{S}_{1:t}=\bm{s}$ lies in a high-dimensional space, it is generally impossible to obtain tight confidence bounds without additional structural assumptions. Specifically, nonparametric estimators suffer from the curse of dimensionality, causing inflated confidence bands and consequently large $\widehat{D}_{t}$. This reflects a fundamental limitation of nonparametric inference—without further assumptions, nontrivial guarantees cannot be achieved in the worst case. To address this issue, one may collect substantially more data or impose structural restrictions that effectively reduce the intrinsic dimension, such as additivity (Meier et al.,, 2009), single-index models (Ichimura,, 1993), or shape constraints (Groeneboom and Jongbloed,, 2014; Chetverikov and Wilhelm,, 2017).
Appendix C Proof of Lemma 4.1
Lemma C.1.
@@
Appendix D Proof of Lemma 4.2
@@
$\mathbb{E}\left[\varepsilon_{t}\mid\mathcal{G}_{t-1}\right]=0.$ (D.2)
@@
$A_{t}\in\sigma\left(\bm{S}_{1:t},\bm{Y}_{1:t-1},\eta_{1:t-1},U_{1:t}\right).$ (D.3)
@@
and Equation (D.1) holds. For Equation (D.2) to hold, we have
@@
where equality $(i)$ holds from Equation (D.3) and equality $(ii)$ follows from $\bm{Y}_{t}$ is independent of $U_{1:t}$. Applying Assumption 4.1 with the above display, it follows that
@@
It is then straightforward to see that Equation (D.2) holds by the definition of $\varepsilon_{t}$. We then conclude that $\{\varepsilon_{t}\}_{t=1}^{T}$ is a martingale difference sequence with respect to $\{\mathcal{G}_{t}\}_{t=1}^{T}$
@@
By Azuma-Hoeffding inequality (see Corollary 2.20 in Wainwright,, 2019), we conclude that $\{\varepsilon_{t}\}_{t=1}^{T}$ is a martingale difference sequence with sub-Gaussian parameter $\sigma^{2}_{\varepsilon}\leq 4$. ∎
Appendix E Proof of Theorem 4.1
@@
Before presenting the proof, we briefly outline the main steps. We first assume that for all $t\in[T]$, $\bm{\theta}^{\star}\in\texttt{BALL}_{t-1}$, where $\texttt{BALL}_{t-1}$ is the confidence ball at step t1t-1 defined in Equation (3.5). Under this assumption, we show that the regret at each step $t\in[T]$ can be decomposed into two components (see Equation (E.9)): the first reflecting the “width” of the confidence ellipsoid in the direction of the chosen decision $\widehat{\bm{\phi}}_{t,A_{t}}$, and the second capturing the imputation error. The former is bounded in Lemma E.1, while the latter is controlled by Equation (E.8). Finally, we select an appropriate sequence $\{\gamma_{t}\}_{t\in[T]}$ to guarantee that $\bm{\theta}^{\star}\in\texttt{BALL}_{t}$ with high probability.
@@
$\bar{\bm{\theta}}_{t}^{\top}\widehat{\bm{\phi}}_{t,A_{t}}$ $=\max_{\bm{\theta}\in\texttt{BALL}_{t-1}}\bm{\theta}^{\top}\widehat{\bm{\phi}}_{t,A_{t}}=\max_{a\in\mathcal{A}}\max_{\bm{\theta}\in\texttt{BALL}_{t-1}}\bm{\theta}^{\top}\widehat{\bm{\phi}}_{t,a}$ (E.1)
where the last equality follows from the way we choose $A_{t}$ as defined in Equation (3.2). Recall that $A^{\star}_{t}$ is the optimal action given by Equation (2.4). The right-hand side of Equation (E.1) is lower bounded by
@@
Taking the above display into Equation (E.1) gives
@@
$\bm{\theta}^{\star\top}\mathbb{E}\left[\bm{\Phi}(\bm{Y}_{t},A^{\star}_{t})\mid\bm{S}_{1:t},A^{\star}_{t}\right]\leq\bar{\bm{\theta}}_{t}^{\top}\widehat{\bm{\phi}}_{t,A_{t}}-\bm{\theta}^{\star\top}\left(\widehat{\bm{\phi}}_{t,A^{\star}_{t}}-\mathbb{E}\left[\bm{\Phi}(\bm{Y}_{t},A^{\star}_{t})\mid\bm{S}_{1:t},A^{\star}_{t}\right]\right).$ (E.2)
@@
$\mathrm{reg}_{t}$ $=\mathbb{E}\left[R(t,A^{\star}_{t})-R(t,A_{t})\mid\bm{S}_{1:t}\right]$ (E.3)
@@
where inequality $(i)$ follows from Equation (E.2) and equality $(ii)$ follows from adding and subtracting $\widehat{\bm{\theta}}_{t}^{\top}\widehat{\bm{\phi}}_{t,A_{t}}$, where $\widehat{\bm{\theta}}_{t}$ is defined in Equation (3.4).
@@
$\left|(\bm{\theta}-\widehat{\bm{\theta}}_{t})^{\top}\bm{\phi}\right|\leq\sqrt{\gamma_{t}\bm{\phi}^{\top}\bm{\Sigma}_{t}^{-1}\bm{\phi}}.$ (E.4)

### v3

@@ 6 NUMERICAL EXPERIMENTS
In this section, we validate our theory and algorithm with simulations on synthetic data and the real Taobao Ad Display/Click dataset (Alibaba, 2018). To facilitate reproducibility, the complete source code is publicly available at https://github.com/Heyan-Zhang/PULSE-UCB.
@@ 6.1 Synthetic Experiments
PULSE-UCB consists of two phases. In pretraining, a context transition model is learned from $N=1000$ historical time series of length $T_{0}=100$ to predict the latent context $W_{t}$. In the online evaluation, the agent runs for $T=1000$ steps. We compare against two benchmarks: (i) OFUL, a naive agent that ignores $W_{t}$ and uses only $\bm{S}_{t}$; (ii) OFUL-Full, an idealized agent with access to the full context $\bm{Y}_{t}=(\bm{S}_{t},W_{t})$. The cumulative regret, averaged over 30 independent trials, is shown in Figure 1. As expected, OFUL-Full achieves the lowest regret since it observes the full context, while OFUL performs worst by ignoring the missing component. In both the linear and nonlinear settings for the missing context, PULSE-UCB performs nearly as well as OFUL-Full, demonstrating the clear benefit of leveraging a predictive model for the unobserved context. To further demonstrate the performance of our approach, we also compared PULSE-UCB with CLBBF (Kim et al., 2023). However, because CLBBF intrinsically relies on a Missing-At-Random (MAR) assumption, it struggles to handle the more general Missing-Not-At-Random (MNAR) structures present in our setting. Consequently, PULSE-UCB consistently outperforms CLBBF and achieves near-oracle performance. Additional experimental details, including sensitivity analyses on pre-training data scales ($N$ and $T_{0}$), are deferred to Appendix M.1.
@@ 6.2 Real-World Experiments
To evaluate our method in a practical setting, we use the public Taobao Ad Display/Click dataset (Alibaba, 2018), which contains 186,730 advertisement display/click records from Taobao.com. Each record includes 83 features describing user and ad attributes such as gender, age, consumption grade, brand, and category. We embed the features into a 32-dimensional space and partition them into 16 observed features ($\bm{S}_{t}$) and 16 unobserved features (${\bm{W}}_{t}$). All algorithms are evaluated on $80\%$ of the data. For PULSE-UCB, we additionally use the remaining $20\%$ for pretraining the context transition model. The action corresponds to selecting an ad (adgroup ID), and the reward is the binary click feedback (1 if clicked, 0 otherwise). Further preprocessing details are deferred to the Appendix N.
We compare PULSE-UCB with three baselines: OFUL, which ignores the missing context; OFUL-Full, which has access to the full context; and CLBBF (Kim et al., 2023), designed for bandits with stochastically missing features. We compare these algorithms over $T\approx 1.5\times 10^{5}$ steps, with $K=20$ arms per step, averaging results over 5 runs. Figure 2 shows that PULSE-UCB greatly outperforms OFUL, highlighting the benefit of context reconstruction, and also surpasses CLBBF, whose mechanism struggles under structural missingness. Notably, PULSE-UCB achieves performance nearly indistinguishable from the ideal OFUL-Full, indicating that the pretraining step not only imputes the missing context but also produces a feature representation well suited for linear bandit learning.
@@
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model.
      Yes: Our problem formulation, theoretical assumptions, and the proposed PULSE-UCB algorithm are detailed in Sections 2 and 3.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm.
      Yes: We provide a theoretical analysis of the regret bound in Section 4 and 5.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries.
      Yes: We provide anonymized source code for all our experiments in the supplementary material.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results.
      Yes: All assumptions for our theoretical results are formally stated in Sections 2-5.
    2. (b)
      Complete proofs of all theoretical results.
      Yes: Complete proofs for all theorems and lemmas are provided in Appendix.
    3. (c)
      Clear explanations of any assumptions.
      Yes: We have provided intuitions and discussions for all our key assumptions.
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL).
      Yes: The code and instructions to reproduce all experimental results are in the supplementary material. For the real-world experiment, our code is based on the public Taobao dataset.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen).
      Yes: We provide all training details, including data splits and hyperparameter settings, in the paper (Section 6.1 and 6.2) as well as the Appendix.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times).
      Yes: All our results are over multiple independent runs and shown with error bars. All shaded areas denote $\pm$ one standard error over multiple runs.
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider).
      No: Our experiments are computationally inexpensive and were run on a standard desktop computer; they do not require any specialized computing infrastructure.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets.
      Yes: We cite the source of the Taobao dataset in Section 6.2.
    2. (b)
      The license information of the assets, if applicable.
      Yes: This Taobao dataset is provided by Alibaba and distributed under the CC BY-NC-SA 4.0 license.
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable.
      Not Applicable: We use an existing public dataset and do not release new assets.
    4. (d)
      Information about consent from data providers/curators.
      Not Applicable: We use a well-established public dataset that has been anonymized.
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content.
      Not Applicable: The dataset is anonymized and does not contain personally identifiable or offensive content.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots.
      Not Applicable: Our research does not involve human subjects or crowdsourcing.
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable.
      Not Applicable: Our research does not involve human subjects.
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation.
      Not Applicable: Our research does not involve human subjects.
@@ Appendix A Additional literature review
AI-assisted decision-making. Recently, there has been growing interest in applying AI, including foundation models, to enhance decision-making. For example, Tianhui Cai et al. (2024); Zhang et al. (2025) design Thompson sampling algorithms for online bandits that treat uncertainty as missing future outcomes, imputing them with pretrained generative models to optimize policies. Chen et al. (2021); Janner et al. (2021) recast offline reinforcement learning as sequence modeling over trajectories to improve decisions, while Lin et al. (2023); Lee et al. (2023) study in-context reinforcement learning, showing that supervised pretraining on past trajectories enables models to approximate algorithms such as LinUCB and Thompson sampling with regret guarantees. Applications include LLM-assisted adaptive experimentation for content delivery (Ye et al., 2025) and human–AI collaboration in linear bandits with resource constraints for healthcare (Cao et al., 2024). Our work focuses on the missing-context issue, specifically in online contextual bandits, and provides near-optimal regret guarantees that guide the use of pretrained models for imputation in this setting.
@@
Imputation in statistics and ML. Imputation has long been a central strategy across statistics and machine learning for handling missing information. A classical example is the EM algorithm (Dempster et al., 1977), which provides a likelihood-based framework for parameter estimation with incomplete data and remains highly influential in this area. In causal inference, imputation is widely used for estimating potential outcomes under counterfactual interventions (Little and Rubin, 2019). From a statistical learning perspective, recent work has incorporated modern machine learning models to deal with missing responses: Xia and Wainwright (2024) propose surrogate training that leverages helper covariates to impute pseudo-responses for unlabeled data, yielding prediction improvements with excess risk guarantees, while Angelopoulos et al. (2023) demonstrate that combining a small labeled set with imputed outcomes enables valid confidence intervals and hypothesis tests. Our work contributes to this line of research by extending imputation-based methods to sequential decision-making with partially observed contexts, and quantify the impact of imputation quality on online learning performance.
Appendix B Practical Motivations for Asymmetric Context Availability
A core premise of our framework is that the auxiliary offline dataset $\mathcal{D}$ contains richer or less noisy contextual information ($Y_{t}$) than what can be feasibly observed during real-time online decision-making ($S_{t}$). This asymmetric context availability is driven by how data are collected in practice. In many real-world environments, online intervention systems must operate under strict privacy constraints, tight computational limits, or high risks of missing data. Conversely, offline datasets are often collected from prior clinical trials, extensive observational cohorts, or archival systems, affording to store and process much richer state variables.
Rather than assuming the online agent can access a complete context capturing all factors that affect the reward, our framework aims for context enrichment: leveraging the rich contextual variables in the auxiliary data to build a more accurate reward approximation offline, which in turn improves the decisions made online using the limited context. We illustrate this setting with two concrete domains:
Digital Health.
In mobile health (mHealth) and digital interventions, multiple sources of context-rich offline data are common, such as prior trials, earlier interventions, and observational cohorts equipped with detailed baseline surveys or high-burden ecological momentary assessments (Liao et al., 2020; Coughlin et al., 2024). By contrast, online intervention systems running on local devices face tight bandwidth limits, battery constraints, and strict privacy regulations. Furthermore, user disengagement and frequent sensor failures often lead to missing or coarsened context in practice (Gonzalez et al., 2025). In our setting, the offline full state $Y_{t}$ represents these rich baseline variables, low-frequency assessments (e.g., stress assessed via validated multi-item scales), and high-resolution wearable signals (e.g., heart-rate variability). The online observed state $S_{t}$, however, is typically restricted to coarser, continuously available signals like step counts or simple self-reports.
Online Education.
In intelligent tutoring systems and online education platforms, the online bandit algorithm typically only observes a limited set of immediate signals $S_{t}$, such as short performance summaries or recent clickstreams. However, educational platforms maintain rich archival databases containing concept-level mastery estimates, comprehensive exam scores, and detailed background information from prior assessments (Lan and Baraniuk, 2016). For instance, in large-scale MOOCs, frameworks like MOOClet utilize archival data (e.g., full homework records and fine-grained logs) to model detailed student states and design adaptive components (Williams et al., 2014). Within our framework, these archival records constitute the auxiliary data $\mathcal{D}$, where $Y_{t}$ aggregates the richer, knowledge-related quantities inferred offline, while the online agent relies solely on the high-frequency, easily observable signals $S_{t}$.
Robustness to Distribution Shift.
While our primary theoretical analysis focuses on the matched-distribution regime where the auxiliary dataset and the online environment share the same context distribution, the learned latent structure often generalizes robustly when this assumption is relaxed. As demonstrated empirically in Appendix M.4, our method maintains strong performance even when the online environment experiences significant covariate shifts relative to the offline auxiliary data. A theoretical justification for this robustness, showing that the rates in Propositions 4.1 and 4.2 are preserved under mild covariate shift, is given in Appendix H.3.
Appendix C Discussion of Assumption 4.1 and Autoregressive Dependence
Assumption 4.1 requires that $\mathbb{E}[\bm{\Phi}(\bm{Y}_{t},a)\mid\bm{Y}_{1:t-1},\eta_{1:t-1},\bm{S}_{1:t}]=\mathbb{E}[\bm{\Phi}(\bm{Y}_{t},a)\mid\bm{S}_{1:t}]$. As noted in the main text, this is satisfied whenever $\bm{Y}_{t}$ is generated as an arbitrary function of $\bm{S}_{1:t}$ together with independent randomness, because the fresh noise at time $t$ carries no information about the past beyond what is already captured by $\bm{S}_{1:t}$. This covers a broad class of models, including all i.i.d. contextual bandit settings as a special case.
The assumption fails, however, when the noise driving $\bm{Y}_{t}$ is itself temporally dependent, as in autoregressive (AR) models. In such cases, past noise innovations $\xi_{1:t-1}$ influence both $\bm{Y}_{1:t-1}$ and the current $\bm{Y}_{t}$, so that conditioning on $\bm{S}_{1:t}$ alone no longer screens off the dependence on the past. The residual $\varepsilon_{t}$ defined in Equation (4.5) then fails to form a martingale difference sequence, and the self-normalized concentration machinery underlying our regret analysis breaks down. Handling $\varepsilon_{t}$ in this regime requires fundamentally different tools. Moreover, existing works on AR bandits suggest that the appropriate techniques depend heavily on the specific form of AR structure, making it difficult to develop a unified treatment. We illustrate this point with two representative examples below.
First, suppose we ignore the observed context $\bm{S}_{1:t}$ and let the latent state evolve via action-dependent dynamics
$W_{t+1}=\mu(a_{t})+\rho(a_{t})\,W_{t}+\xi_{t},$
where $\xi_{t}$ is i.i.d. noise. The reward satisfies $R(t,a_{t})=\bm{\theta}^{\star\top}\bm{\Phi}(\bm{Y}_{t},a_{t})+\widetilde{\eta}_{t}$, giving a model whose transition coefficients depend on the agent’s action. This is the action-driven AR bandit studied by Bacchiocchi et al. (2024), for which modified UCB-type algorithms exploit the action-dependent structure to achieve the optimal $\mathcal{O}(\sqrt{T})$ regret. Crucially, however, this structure is absent in our problem: the auxiliary historical data contain no action information, and the contextual process evolves according to fixed, action-independent dynamics. Techniques designed for action-driven AR dependence therefore do not apply.
Second, consider a form of AR dependence more representative of our setting. Let
$W_{t}=\alpha\,W_{t-1}+\xi_{t},\qquad\xi_{t}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,1),\quad|\alpha|<1,$
and set $\bm{Y}_{t}=W_{t}$ with feature map $\bm{\Phi}(\bm{Y}_{t},a)=\beta_{a}W_{t}$. The reward becomes $R(t,a)=\bm{\theta}^{\star}\,\beta_{a}\,W_{t}+\eta_{t}$. Here the AR dependence resides in a latent process that influences the reward only multiplicatively through the action, producing a fundamentally different statistical structure from the first example. This setting corresponds to the latent autoregressive bandit model of Trella et al. (2024), where the analysis requires Kalman-filter-type estimators and strong linear-Gaussian assumptions, and the resulting regret may exceed $\mathcal{O}(\sqrt{T})$.
These two examples illustrate that it is currently not straightforward to propose a unified approach for handling AR dependence in bandit problems. The difficulty of the learning problem hinges critically on how autoregressive dependence enters the model, and seemingly minor modeling choices, such as whether the AR dependence acts on the observed reward or on a latent state, and whether it is action-driven or governed by fixed dynamics, lead to qualitatively different algorithms and regret guarantees. Furthermore, both examples above assume that the feature map $\bm{\Phi}$ is linear in the contextual variables; once nonlinear dependencies are introduced, it is unclear how to adapt existing techniques or whether the same regret rates remain achievable.
Taken together, these observations underscore that extending our framework to accommodate autoregressive noise would require a case-by-case analysis tied to the specific AR structure at hand, rather than a single unified extension of Assumption 4.1. We therefore leave this direction to future work.
Appendix D A data-driven approach for choosing $D_{t}$
@@
    $\big|\mu_{p}(\bm{s})-\mu_{\widehat{p}}(\bm{s})\big|\;\leq\;\underbrace{\big|\mu_{p}(\bm{s})-\mu_{\widehat{p}_{0}}(\bm{s})\big|}_{\text{controlled by the UCB}}\;+\;\underbrace{\big|\mu_{\widehat{p}_{0}}(\bm{s})-\mu_{\widehat{p}}(\bm{s})\big|}_{\text{fully data-computable}}.$ (D.1)
@@
$\mathcal{E}_{t}(p,\widehat{p};\bm{s})\;\leq\;\sup_{\bm{s}\in\mathcal{G}_{t}}r_{1-\alpha}(\bm{s})\;+\;\big|\mu_{\widehat{p}_{0}}(\bm{s})-\mu_{\widehat{p}}(\bm{s})\big|,$ (D.2)
@@
    UCB machinery for the reference fit $\widehat{p}_{0}$. Two widely used choices are: (i) local-polynomial regression with robust bias correction (RBC), whose studentized process admits valid simultaneous bands and is robust to MSE-optimal bandwidth choice; the quantiles are obtained via Gaussian/multiplier bootstrap of the sup-statistic; (ii) penalized splines with simultaneous bands via volume-of-tube or bootstrap calibrations. [^3] See, e.g., Calonico et al. (2018, 2022) for RBC-based bands and Krivobokova et al. (2010) for spline bands; multiplier bootstrap for suprema is treated in Chernozhukov et al. (2014).
@@
When the observed context $\bm{S}_{1:t}=\bm{s}$ lies in a high-dimensional space, it is generally impossible to obtain tight confidence bounds without additional structural assumptions. Specifically, nonparametric estimators suffer from the curse of dimensionality, causing inflated confidence bands and consequently large $\widehat{D}_{t}$. This reflects a fundamental limitation of nonparametric inference—without further assumptions, nontrivial guarantees cannot be achieved in the worst case. To address this issue, one may collect substantially more data or impose structural restrictions that effectively reduce the intrinsic dimension, such as additivity (Meier et al., 2009), single-index models (Ichimura, 1993), or shape constraints (Groeneboom and Jongbloed, 2014; Chetverikov and Wilhelm, 2017).
Appendix E Proof of Lemma 4.1
Lemma E.1.
@@
Appendix F Proof of Lemma 4.2
@@
$\mathbb{E}\left[\varepsilon_{t}\mid\mathcal{G}_{t-1}\right]=0.$ (F.2)
@@
$A_{t}\in\sigma\left(\bm{S}_{1:t},\bm{Y}_{1:t-1},\eta_{1:t-1},U_{1:t}\right).$ (F.3)
@@
and Equation (F.1) holds. For Equation (F.2) to hold, we have
@@
where equality $(i)$ holds from Equation (F.3) and equality $(ii)$ follows from $\bm{Y}_{t}$ is independent of $U_{1:t}$. Applying Assumption 4.1 with the above display, it follows that
@@
It is then straightforward to see that Equation (F.2) holds by the definition of $\varepsilon_{t}$. We then conclude that $\{\varepsilon_{t}\}_{t=1}^{T}$ is a martingale difference sequence with respect to $\{\mathcal{G}_{t}\}_{t=1}^{T}$
@@
By Azuma-Hoeffding inequality (see Corollary 2.20 in Wainwright, 2019), we conclude that $\{\varepsilon_{t}\}_{t=1}^{T}$ is a martingale difference sequence with sub-Gaussian parameter $\sigma^{2}_{\varepsilon}\leq 4$. ∎
Appendix G Proof of Theorem 4.1
@@
Before presenting the proof, we briefly outline the main steps. We first assume that for all $t\in[T]$, $\bm{\theta}^{\star}\in\texttt{BALL}_{t-1}$, where $\texttt{BALL}_{t-1}$ is the confidence ball at step t1t-1 defined in Equation (3.5). Under this assumption, we show that the regret at each step $t\in[T]$ can be decomposed into two components (see Equation (G.9)): the first reflecting the “width” of the confidence ellipsoid in the direction of the chosen decision $\widehat{\bm{\phi}}_{t,A_{t}}$, and the second capturing the imputation error. The former is bounded in Lemma G.1, while the latter is controlled by Equation (G.8). Finally, we select an appropriate sequence $\{\gamma_{t}\}_{t\in[T]}$ to guarantee that $\bm{\theta}^{\star}\in\texttt{BALL}_{t}$ with high probability.
@@
$\bar{\bm{\theta}}_{t}^{\top}\widehat{\bm{\phi}}_{t,A_{t}}$ $=\max_{\bm{\theta}\in\texttt{BALL}_{t-1}}\bm{\theta}^{\top}\widehat{\bm{\phi}}_{t,A_{t}}=\max_{a\in\mathcal{A}}\max_{\bm{\theta}\in\texttt{BALL}_{t-1}}\bm{\theta}^{\top}\widehat{\bm{\phi}}_{t,a}$ (G.1)
where the last equality follows from the way we choose $A_{t}$ as defined in Equation (3.2). Recall that $A^{\star}_{t}$ is the optimal action given by Equation (2.4). The right-hand side of Equation (G.1) is lower bounded by
@@
Taking the above display into Equation (G.1) gives
@@
$\bm{\theta}^{\star\top}\mathbb{E}\left[\bm{\Phi}(\bm{Y}_{t},A^{\star}_{t})\mid\bm{S}_{1:t},A^{\star}_{t}\right]\leq\bar{\bm{\theta}}_{t}^{\top}\widehat{\bm{\phi}}_{t,A_{t}}-\bm{\theta}^{\star\top}\left(\widehat{\bm{\phi}}_{t,A^{\star}_{t}}-\mathbb{E}\left[\bm{\Phi}(\bm{Y}_{t},A^{\star}_{t})\mid\bm{S}_{1:t},A^{\star}_{t}\right]\right).$ (G.2)
@@
$\mathrm{reg}_{t}$ $=\mathbb{E}\left[R(t,A^{\star}_{t})-R(t,A_{t})\mid\bm{S}_{1:t}\right]$ (G.3)
@@
where inequality $(i)$ follows from Equation (G.2) and equality $(ii)$ follows from adding and subtracting $\widehat{\bm{\theta}}_{t}^{\top}\widehat{\bm{\phi}}_{t,A_{t}}$, where $\widehat{\bm{\theta}}_{t}$ is defined in Equation (3.4).
@@
$\left|(\bm{\theta}-\widehat{\bm{\theta}}_{t})^{\top}\bm{\phi}\right|\leq\sqrt{\gamma_{t}\bm{\phi}^{\top}\bm{\Sigma}_{t}^{-1}\bm{\phi}}.$ (G.4)